calculate the price elasticity of demand for a commodity when its price increase by 50% and quantity demanded falls to 240 units from 300 units


1
Expert's answer
2021-01-12T07:37:52-0500

Price elasticity of demand "=\%" change in quantity"\/" "\%" change in price

"\%" change in price "=50\%"

"\%" change in quantity "=\frac{(240-300)}{300}\times 100\%=-20\%"

Price elasticity of demand "=-20\/50=-0.4"

Price elasticity of demand"=|-0.4|=0.4"